The pressure in a vessel that contained pure oxygen dropped from 1000mm to 750mm in47 minutes as the oxygen is leaked through a small hole.when the same vessel was filled with another gas the pressure dropped from 1000mm to 750 in 74 minutes.What is the molecular weight of the second gas?

Grade:11

1 Answers

Vikas TU
7 years ago
Rate of diffusion,
dP/dt => (1000-750)/47 = PA/root(32*T)..........(1)
 (1000-750)/74 = PA/root(M*T).............(2)
dividing both the eqns. we get,
M = (74/47)^2*32 = > 79.32 approx.
that is the molar mass of the Br gas.